Shift summary - Eve 11/09/99

Julie Whitmore (jaws@fnal.gov)
Wed, 10 Nov 1999 00:08:14 -0600 (CST)

Shift Summary for Eve shift, 09/Nov/99

On Shift:Ricardo, Julie

Hours of beam available at all intensities: 8.25
Hours of beam available at nominal intensity: 8.25
Hours of E799 data (special runs not included): 7.25
Hours of special runs (laser scans, muon runs...): 0.0

Accounting of lost time: Run aborted due to dirty ktev4 drive and/or
bad tape. Ended run early to update CSQ6 and DCT0
calibration constants.

========================================================================
Runs ended this shift:
Run 14933 - 23.8M events - BEAM_799
Run 14936 - 11.7M events - BEAM_799

Tapes brought to FCC this shift:
(including tapes for online split)

UP*115-123,UPS139-40

UPM022-024,UPF027-030,UPH016-018,UPK014-015,UPT009,UPX009,UPE008
(vaulted by Arun)

========================================================================
Accesses during shift: None.

========================================================================
Overview of shift:
Relatively smooth data taking interrupted by Taku spontaneously aborting
a run. Beam structure is variable (sometimes 2Hz, sometimes smooth).
Updated CSQ6 and DCT0 constants.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14933 15:30 Arrive to find Margaret Votava working on
DOT problem. Flying blind w/o tape usage
information.

15:45 Phone MCR to complain about 2Hz structure.
No obvious improvment made.

16:15 Alcohol bubbler filled by DN.

16:10 Margaret gets DOT working (see logbook XXII,
p.201 for details). DOT struggles to catch
up to run.

16:23 Tapes over 100% full. Taku announces he is
full. We believed the run ended gracefully,
but CsI and DC experts find that calibration
files are missing (DCT0 and CSQ6) for that
run.

14934 16:30 Start run 14934.

16:50 Run aborts. ktev4 tape drive 1 needs
cleaning. Page Sydney. Margaret sees
"hardware error" in log file. Decide to
clean drive and use new tape.

14936 17:35 Decide to overwrite tapes. Run init hangs
at stage 2. First dart fresh start fails
because of DAMP_FWDR problem. Second fresh
start works. Start run 14936. CsI 1096 gain
drift is down to -11%.

17:50 HyperCP folks phone to ask about beam
structure, which has been pretty bad during
our entire shift.

33-37 18:20 Beam gone.

48 18:35 Beam smooths out. We're wondering if HyperCP
complained too.

58-60 18:50 Missing and partial spills.

19:00 Onsplit tape mount - UPH019.

19:25 Onsplit tape mount - UPF031.

21:50 Onsplit tape mount - UPM025.

Discussed the gain drift and E/P situation
with BobT and Ed. Decide to end run early
and update CSQ6 constants. Ashkan wants
DCT0 updated also to fix ineff DC wire
(PL7,W123).

23:10 Stop run. Page Ashkan. Update CSQ6 and
DCT0 constants.

23:15 Onsplit tape mount - UPLA06.

14937 23:20 Start run. Update laser gain reference
plot. This will be picked up at the
beginning of the next run.

23:30 Phone MCR to complain yet again about the
beam structure. They are working on it.

========================================================================
Follow up Notes on Previous Shift
Install CSQ6 constants and update laser gain reference plot. This should
rectify 1096 gain drift problem.

========================================================================
Items to Follow up on Next Shift
Updated laser gain reference plot. Should update at the start of the next
run.

Noticed L3out/L3in has been slowly dropping (down to 0.135). Any ideas?